An object moving on the x axis with a con- stant acceleration increases its x coordinate by 232 m in a time of 2.69 s and has a velocity of 20 m/s at the end of this time.
Determine the acceleration of the object
1
Expert's answer
2020-02-27T10:04:52-0500

The distance traveled by the object


"d=\frac{v_i+v_f}{2}t,"

so, the initial velocity


"v_i=\frac{2d}{t}-v_f=\frac{2\times 232}{2.69}-20=152.5\:\rm m/s."

The acceleration of the object


"a=\frac{v_f-v_i}{t}=\frac{20-152.5}{2.69}=-49.25\:\rm m/s^2."